# Integrating Multi-Preconditioned Conjugate Gradient with Additive Multigrid Strategy

### v1

@@ 1 Problem definition
To numerically solve (1), the finite element methods (FEM) are utilized for discretization. This approach effectively transforms the continuous diffusion problem into an algebraic system of equations, typically formulated as $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{b}$.
@@ 2.1 Additive multigrid method
The multigrid method we usually employ in practice is a multiplicative multigrid method. As in the $V$/$W$-cycle, the global correction is computed using the successive application of the smoothing steps on each level while going first from the fine-to-coarse levels and then from the coarse-to-fine levels 7_hackbusch_multi-grid_1986 . In contrast, that additive MG method allows for the application of the smoothing operations in parallel, completely independent of each other 2_bastian_additive_1998 . Thus, the additive MG method presents a benefit over traditional multiplicative MG techniques due to its inherently parallel nature, which can be advantageous in modern high-performance computing environments 5_chen_multigrid_2001 ; 6_darwish_parallelization_2008 .
@@
To realize a multigrid method, we assume that the multilevel hierarchy of FE spaces is constructed either in a geometric manner, i.e., by constructing a hierarchy of nested meshes, or in an algebraic manner, i.e., by employing some agglomeration methods. We denote this hierarchy of nested FE spaces as ${\pazocal{V}_{0}\subset\pazocal{V}_{1}\subset\cdots\subset\pazocal{V}_{L}}$, where $L$ denotes the finest level and 00 denotes the coarsest level of the multilevel hierarchy. On this hierarchy of FE spaces, we construct the prolongation operators that can transfer the functions from a coarse level to the finest level, given as
$P_{\ell}^{L}:\pazocal{V}_{\ell}\to\pazocal{V}_{L},\qquad\forall\ell\in\{0,1,\ldots,L\}.$
Similarly, we also define the restriction operators that transfer the residuals from the finest level to a coarse level, given as
$R_{L}^{\ell}:\pazocal{V}^{\ast}_{L}\to\pazocal{V}^{\ast}_{\ell},\qquad\forall\ell\in\{0,1,\ldots,L\},$
where for the linear MG method, the restriction operators can be constructed as the transpose of the prolongation operator, e.g., $R_{L}^{\ell}=(P_{\ell}^{L})^{\top}$. Also, for $\ell=L$, the transfer operators degenerate to the identity operator. We note that these transfer operators for the additive MG method can be constructed explicitly or computed as composite operators by successively applying the transfer operators defined between two consecutive levels. Additionally, we define smoothing operators on each level of the multilevel hierarchy except the coarsest level, given as:
$\pazocal{S}_{i}:\pazocal{V}_{\ell}^{\ast}\to\pazocal{V}_{\ell},\qquad\forall\ell\in\{1,2,\ldots,L\},$
@@
while on the coarsest level, we define an approximate inverse of the coarse level operator $\pazocal{A}_{0}$ as $\pazocal{S}_{0}:\pazocal{V}^{\ast}_{0}\to\pazocal{V}_{0}$. As a single step of the smoothing iteration may not be sufficient to smooth the error associated with the high-frequency components of the error, it becomes essential to perform multiple smoothing iterations, which we define as $\nu$. With these operators, we can define the update of the additive multigrid method as
$\boldsymbol{x}_{k+1}=\boldsymbol{x}_{k}+\sum_{\ell=0}^{L}P_{\ell}^{L}\pazocal{S}^{\nu}_{\ell}R_{L}^{\ell}\boldsymbol{r}_{k}\qquad\text{where, }\boldsymbol{r}_{k}=\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{k}.$ (2)
@@
As we can observe, in (2) the corrections from all levels are equally weighted, and the method might also diverge if the corrections are not appropriately damped. Due to this reason, the additive MG method is usually employed as a preconditioner, where the optimal damping parameter is computed by the outer iterative method, usually a Krylov subspace method.
@@
In practice, the additive MG method has cheaper synchronization costs than the standard multiplicative MG method on the parallel architectures 1_bastian_load_1998 . However, the convergence of the additive MG-PCG method remains slower than its multiplicative counterpart, which prohibits the usage of the additive MG-PCG method as a solution method for large-scale problems. We aim to improve the convergence of the overall methodology by combining the additive MG method and the multipreconditioned CG method.
2.2 Multipreconditioned Conjugate Gradient method
@@
$\boldsymbol{P}_{k}=[P^{L}_{0}S_{0}R_{L}^{0}\boldsymbol{r}_{k}\ |\ P^{L}_{1}S_{1}^{\nu}R_{L}^{1}\boldsymbol{r}_{k}\ |\cdots\ |\ P^{L}_{L}S^{\nu}_{L}R_{L}^{L}\boldsymbol{r}_{k}]\in\mathbb{R}^{n\times(L+1)}.$
@@
$\mathbb{R}^{L+1}\ni\boldsymbol{\alpha}=(\boldsymbol{P}_{k}^{\top}\boldsymbol{A}\boldsymbol{P}_{k})^{-1}(\boldsymbol{P}_{k}\boldsymbol{r}_{k}),$
@@
$\boldsymbol{p}_{k+1}=\sum_{\ell=0}^{L}\alpha_{\ell}P_{\ell}^{L}\pazocal{S}^{\nu}_{\ell}R_{L}^{\ell}\boldsymbol{r}_{k}.$ (3)
With this approach, we can repurpose the smoothing operations on the different levels of the multilevel hierarchy as actions of different preconditioners. This approach leverages different subspaces to reduce the error in different frequency spectrum, thereby constructing a more robust preconditioning approach. In the context of the MPCG method, it is crucial to select preconditioners with distinct properties, such that the preconditioned residuals do not become linearly dependent; otherwise, the matrix $\boldsymbol{P}_{k}^{\top}\boldsymbol{A}\boldsymbol{P}_{k}$ can become non-singular. If the FE spaces in the multilevel hierarchy are $H_{1}$-orthogonal to each other, we can significantly enhance the search space, leading to improved convergence and also avoid the matrix $\boldsymbol{P}_{k}^{\top}\boldsymbol{A}\boldsymbol{P}_{k}$ from becoming singular. In Algorithm 1, we can see the detailed algorithm for combining the additive MG and the MPCG method.
@@
  normal-$\triangleright$ Compute residual
$\boldsymbol{z}^{\ell}_{0}\mapsfrom P_{\ell}^{L}\pazocal{S}_{\ell}^{\nu}R_{L}^{\ell}\boldsymbol{r}_{0}\quad\forall\ell\in\{0,1,\ldots,L\}$ ;
  normal-$\triangleright$ Corrections from all levels
@@
  normal-$\triangleright$ Preconditioned residuals from $L$ levels
@@ 1 $k\mapsfrom 0$;
       $\boldsymbol{\alpha}\mapsfrom(\boldsymbol{P}_{k}^{\top}\boldsymbol{A}\boldsymbol{P}_{k})^{-1}(\boldsymbol{P}_{k}\boldsymbol{r}_{k})$ ;
        normal-$\triangleright$ Computing $\{\alpha_{\ell}\}_{\ell=0,1,\ldots,L}$
@@
        normal-$\triangleright$ updating the current iterate
@@
        normal-$\triangleright$ updating the residual
       $\boldsymbol{z}^{\ell}_{k+1}\mapsfrom P_{\ell}^{L}\pazocal{S}_{\ell}^{\nu}R_{L}^{\ell}\boldsymbol{r}_{k+1}\quad\forall\ell\in\{0,1,\ldots,L\}$ ;
        normal-$\triangleright$ Corrections from all levels
@@
        normal-$\triangleright$ Preconditioned residuals
@@
        normal-$\triangleright$ Constructing $\beta$
@@ 3 Numerical Experiments
We investigate the performance of the additive-MG multipreconditioned CG method using two numerical examples.
Example 1
Anisotropic diffusion: We consider a domain $\Omega=(0,1)^{2}$ with anisotropic diffusion. The values considered for $k_{xx}$ range across several orders of magnitude, $k_{xx}\in\{10^{n}\}_{n=-7,-6,\ldots,0}$, while we set $k_{yy}=1$. We utilize a mesh with $160\times 160$ elements, which we set as the finest level in the multilevel hierarchy. In the multigrid method, we consider a nested hierarchy of meshes with $4$ levels, while on the coarsest level, we have only $20\times 20$ elements. Within this framework, we employ the symmetric Successive Over-Relaxation (SSOR) method as a smoother and perform $6$ smoothing steps on each level. We note that a line smoother could be employed to further enhance the performance of the multigrid (MG) method, although it is not utilized in the current setup. Furthermore, for the multi-preconditioned CG method, we store $5$ pairs of vectors to estimate the value of the parameter $\beta$.
@@
Example 2
Pressure distribution in the fracture network: In this example, we again consider a domain $\Omega=(0,1)^{2}$. We aim to model the pressure distribution in a fracture network, considering various parameters influencing the flow dynamics 11_nestola_novel_2020 . The fracture network is characterized by a uniform fracture thickness ($\delta$) of $10^{-4}$. The matrix permeability ($k_{m}$), a measure of the ability of the matrix to transmit fluids, is set to $1$. The fracture permeability ($k_{f}$) is chosen as $k_{f}\in\{10^{n}\}_{n=-4,-3,\ldots,4}$, ranging from the blocking to conducting phase. The model involves approximately $3\cdot 10^{5}$ degrees of freedom (DoFs). We enforce $g_{N}=0$ as the Neumann boundary condition on the top and bottom edge of the domain, while $g_{N}=1$ on the left part of the boundary and $u_{D}=1$ is prescribed on the right edge of the domain. A sketch of the fracture network can be seen in Figure 0(a), while the fracture flow in the network for the blocking fractures can be seen in Figure 0(b) and conducting fractures in Figure 0(c). To construct a multilevel hierarchy, we employed an Algebraic Multigrid (AMG) method, utilizing the BoomerAMG from Hypre framework 8_henson_boomeramg_2002 . For the solution methods, we employ the same strategies detailed in Example 1.
@@ 3.1 Comparison with the other solution methods
In this section, we compare the convergence behavior of the additive-MG MPCG method with the additive-MG PCG method and the multiplicative MG ($V(3,3)$) PCG method. For this study, we consider two special test cases for this comparison for Example 1 and Example 2.
@@
For the Example 1, we consider a high anisotropic case ($k_{xx}=10^{-6}$, $k_{yy}=1$) and a low anisotropic case ($k_{xx}=10^{-1}$, $k_{yy}=1$). As shown in Figure 2, in both cases, the MPCG method outperforms the standard PCG method when preconditioned with the additive-MG method. However, as expected, the multiplicative MG-PCG method outperforms both methods. For Example 1, we have employed a multigrid method with geometric hierarchy using the uniform refinement strategy. Hence, we can see that the multilevel hierarchy is not optimal for the highly anisotropic case. Consequently, we observe increased iterations required for convergence across all three solution methods for the highly anisotropic case. For the highly anisotropic case, the performance of the additive MG-MPCG method is closer to the MG($V$)-PCG method.
To understand this behavior in more detail, we carried out the same experiment with more cases with slowly increasing anisotropy. Figure 4 (left) shows that less anisotropic problems converge substantially faster than their highly anisotropic counterparts. Among the tested methods, the additive-MG-PCG requires the highest number of iterations for convergence, and this requirement amplifies more noticeably with increasing anisotropy. Interestingly, the additive-MG-MPCG method, while exhibiting a similar trend, does not experience as significant a rise in number of iterations. Its convergence rate grows at a rate comparable to the MG($V$)-PCG methods.
@@
For Example 2, we consider two different test cases with a varying range of fracture permeabilities, where the two extreme cases are chosen, the blocking fractures ($k_{m}=1,k_{f}=10^{-4}$) and the conducting fractures ($k_{m}=1,k_{f}=10^{4}$). In Figure 3, we observe the convergence of the three solution methods for blocking and conducting cases. Similar to the previous example, the additive-MG-MPCG method outperforms the additive-MG-PCG method for both test cases. In this example, we utilize BoomerAMG to build the multilevel hierarchy, which takes into account the permeabilities of the fractures when constructing the transfer operators. Such agglomeration based construction enhances the quality of the multilevel hierarchy compared to that in Example 1, resulting in a more effective hierarchy of the subspaces. We can observe the effect of the transfer operators in Figure 4 (right). In this case, the number of iterations for additive-MG-PCG and additive-MG-MPCG methods remains relatively constant across different permeabilities, with the additive-MG-MPCG method showing a slight increase in iterations as permeability increases.

### v2

# Integrating Additive Multigrid with Multipreconditioned Conjugate Gradient Method

@@ 1 Problem definition
To numerically solve (1), the finite element method (FEM) is utilized for discretization. This approach effectively transforms the continuous diffusion problem into an algebraic system of equations, typically formulated as $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{b}$.
@@ 2.1 Additive multigrid method
The multigrid method we usually employ in practice is a multiplicative multigrid method. As in the $V$/$W$-cycle, the global correction is computed using the successive application of the smoothing steps on each level while going first from the fine-to-coarse levels and then from the coarse-to-fine levels 7_hackbusch_multi-grid_1986 . In contrast, the additive MG method allows for the application of the smoothing operations in parallel, completely independent of each other 2_bastian_additive_1998 . Thus, the additive MG method presents a benefit over traditional multiplicative MG techniques due to its inherently parallel nature, which can be advantageous in modern high-performance computing environments 5_chen_multigrid_2001 ; 6_darwish_parallelization_2008 .
@@
To realize a multigrid method, we assume that the multilevel hierarchy of FE spaces is constructed either in a geometric manner, i.e., by constructing a hierarchy of nested meshes, or in an algebraic manner, i.e., by employing some agglomeration methods. We denote this hierarchy of nested FE spaces as ${\pazocal{V}_{0}\subset\pazocal{V}_{1}\subset\cdots\subset\pazocal{V}_{L}}$, where $L$ denotes the finest level and 00 denotes the coarsest level of the multilevel hierarchy. On this hierarchy of FE spaces, we construct the prolongation operators that can transfer the functions from a coarse level to the finest level, given as $P_{\ell}^{L}:\pazocal{V}_{\ell}\to\pazocal{V}_{L},\text{for }\ell\in\{0,1,\ldots,L\}.$ Similarly, we also define the restriction operators that transfer the residuals from the finest level to a coarse level, given as $R_{L}^{\ell}:\pazocal{V}^{\ast}_{L}\to\pazocal{V}^{\ast}_{\ell},\text{for }\ell\in\{0,1,\ldots,L\},$ where for the linear MG method, the restriction operators can be constructed as the transpose of the prolongation operator, e.g., $R_{L}^{\ell}=(P_{\ell}^{L})^{\top}$, and $\pazocal{V}^{\ast}_{\ell}$ denotes the dual space of $\pazocal{V}_{\ell}$. Also, for $\ell=L$, the transfer operators degenerate to the identity operator. We note that these transfer operators for the additive MG method can be constructed explicitly or computed as composite operators by successively applying the transfer operators defined between two consecutive levels. Additionally, we define smoothing operators on each level of the multilevel hierarchy except the coarsest level as: $\pazocal{S}_{\ell}:\pazocal{V}^{\ast}_{\ell}\to\pazocal{V}_{\ell},\text{for }\ell\in\{1,2,\ldots,L\},$ where the smoothing iterations for given initial guess $\boldsymbol{c}_{0}$ on level $\ell$ are given as
$\pazocal{S}_{\ell}^{\nu}(R_{L}^{\ell}\boldsymbol{b}):=(\pazocal{I}-\pazocal{M}_{\ell}^{-1}\pazocal{A}_{\ell})^{\nu}\boldsymbol{c}_{0}+\sum_{i=0}^{\nu-1}(\pazocal{I}-\pazocal{M}_{\ell}^{-1}\pazocal{A}_{\ell})^{i}\pazocal{M}_{\ell}^{-1}(R_{L}^{\ell}\boldsymbol{b}),$ (2)
@@
where $\pazocal{M}_{\ell}$ and $\pazocal{I}-\pazocal{M}_{\ell}^{-1}\pazocal{A}_{\ell}$ denote a preconditioner and an iteation matrix induced by the smoother. On the coarsest level, we define an approximate inverse of the coarse level operator $\pazocal{A}_{0}$ as $\pazocal{S}_{0}:\pazocal{V}_{0}^{\ast}\to\pazocal{V}_{0}$. As a single step of the smoothing iteration may not be sufficient to smooth the error associated with the high-frequency components of the error, it becomes essential to perform multiple smoothing iterations, which we define as $\nu$. With these operators, we can define the update of the additive multigrid method as
$\boldsymbol{x}_{k+1}=\boldsymbol{x}_{k}+P_{0}^{L}\pazocal{S}_{0}(R^{0}_{L}\boldsymbol{r}_{k})+\sum_{\ell=1}^{L}P_{\ell}^{L}\pazocal{S}^{\nu}_{\ell}(R^{\ell}_{L}\boldsymbol{r}_{k}),$ (3)
@@
where $\boldsymbol{r}_{k}=\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}_{k}$. As we can observe, in (3) the corrections from all levels are equally weighted, and the method might also diverge if the corrections are not appropriately damped. Due to this reason, the additive MG method is usually employed as a preconditioner, where the optimal damping parameter is computed by the outer iterative method, usually a Krylov subspace method.
@@
In practice, the additive MG method has cheaper synchronization costs than the standard multiplicative MG method on parallel architectures 1_bastian_load_1998 . However, the convergence of the additive MG-PCG method remains slower than its multiplicative counterpart, which prohibits the usage of the additive MG-PCG method as a solution method for large-scale problems. We aim to improve the convergence of the overall methodology by combining the additive MG method and the multipreconditioned CG method.
2.2 Multipreconditioned Conjugate Gradient method (MPCG)
@@
$\boldsymbol{P}_{k}=[P^{L}_{0}\pazocal{S}_{0}(R_{L}^{0}\boldsymbol{r}_{k})\ |\ P^{L}_{1}\pazocal{S}_{1}^{\nu}(R_{L}^{1}\boldsymbol{r}_{k})\ |\cdots\ |\ P^{L}_{L}\pazocal{S}^{\nu}_{L}(R_{L}^{L}\boldsymbol{r}_{k})]\in\mathbb{R}^{n\times(L+1)}.$
@@
$\mathbb{R}^{L+1}\ni\boldsymbol{\alpha}=(\boldsymbol{P}_{k}^{\top}\boldsymbol{A}\boldsymbol{P}_{k})^{-1}(\boldsymbol{P}_{k}^{\top}\boldsymbol{r}_{k}),$
@@
$\boldsymbol{P}_{k}\boldsymbol{\alpha}=\boldsymbol{p}_{k}=\alpha_{0}P_{0}^{L}\pazocal{S}_{0}(R_{L}^{0}\boldsymbol{r}_{k})+\sum_{\ell=1}^{L}\alpha_{\ell}P_{\ell}^{L}\pazocal{S}^{\nu}_{\ell}(R_{L}^{\ell}\boldsymbol{r}_{k}).$ (4)
With this approach, we can repurpose the smoothing operations on the different levels of the multilevel hierarchy as actions of different preconditioners. This approach leverages different subspaces to reduce the error in different frequency of the spectrum, thereby constructing a more robust preconditioning approach. In the context of the MPCG method, it is crucial to select preconditioners with distinct properties, such that the preconditioned residuals do not become linearly dependent; otherwise, the matrix $\boldsymbol{P}_{k}^{\top}\boldsymbol{A}\boldsymbol{P}_{k}$ can become singular. If the FE spaces in the multilevel hierarchy are $H_{1}$-orthogonal to each other, we can significantly enhance the search space, leading to improved convergence and also avoid the matrix $\boldsymbol{P}_{k}^{\top}\boldsymbol{A}\boldsymbol{P}_{k}$ from becoming singular. In Algorithm 1, we can see the detailed algorithm for combining the additive MG and the MPCG method.
@@
  $\triangleright$ Compute residual
$\boldsymbol{z}^{\ell}_{0}\mapsfrom P_{\ell}^{L}\pazocal{S}_{\ell}^{\nu}(R_{L}^{\ell}\boldsymbol{r}_{0})\quad\forall\ell\in\{0,1,\ldots,L\}$ ;
  $\triangleright$ Corrections from all levels
@@
  $\triangleright$ Preconditioned residuals from $L$ levels
@@ 1 $k\mapsfrom 0$;
       $\boldsymbol{\alpha}\mapsfrom(\boldsymbol{P}_{k}^{\top}\boldsymbol{A}\boldsymbol{P}_{k})^{-1}(\boldsymbol{P}_{k}^{\top}\boldsymbol{r}_{k})$ ;
        $\triangleright$ Computing $\{\alpha_{\ell}\}_{\ell=0,1,\ldots,L}$
@@
        $\triangleright$ updating the current iterate
@@
        $\triangleright$ updating the residual
       $\boldsymbol{z}^{\ell}_{k+1}\mapsfrom P_{\ell}^{L}\pazocal{S}_{\ell}^{\nu}(R_{L}^{\ell}\boldsymbol{r}_{k+1})\quad\forall\ell\in\{0,1,\ldots,L\}$ ;
        $\triangleright$ Corrections from all levels
3
@@
        $\triangleright$ Preconditioned residuals
@@
        $\triangleright$ Constructing $\beta$
@@ 3 Numerical Experiments
We investigate the performance of the additive MG-MPCG method using two numerical examples.
@@
Example 1: Anisotropic diffusion: We consider a domain $\Omega=(0,1)^{2}$ with anisotropic diffusion. The values considered for $k_{xx}$ range across several orders of magnitude, $k_{xx}\in\{10^{n}\}_{n=-7,-6,\ldots,0}$, while we set $k_{yy}=1$. We utilize a mesh with $160\times 160$ elements, which we set as the finest level in the multilevel hierarchy. In the multigrid method, we consider a nested hierarchy of meshes with $4$ levels, while on the coarsest level, we have only $20\times 20$ elements.
Example 2: Pressure distribution in the fracture network: In this example, we consider the domain $\Omega=(0,1)^{2}$ to model pressure distribution in a fracture network, considering various flow dynamics parameters 11_nestola_novel_2020 . The fracture network is characterized by a uniform fracture thickness ($\delta$) of $10^{-4}$. The matrix permeability ($k_{m}$), a measure of the ability of the matrix to transmit fluids, is set to $1$. The fracture permeability ($k_{f}$) is chosen as $k_{f}\in\{10^{n}\}_{n=-4,-3,\ldots,4}$, ranging from the blocking to conducting phase. The model involves approximately $3\cdot 10^{5}$ degrees of freedom (DoFs). We enforce $g_{N}=0$ as the Neumann boundary condition on the top and bottom edge of the domain, while $g_{N}=1$ on the left part of the boundary and $u_{D}=1$ is prescribed on the right edge of the domain. A sketch of the fracture network can be seen in Figure 1(a), while the fracture flow in the network for the blocking fractures can be seen in Figure 1(b) and conducting fractures in Figure 1(c). An Algebraic Multigrid (AMG) method using BoomerAMG from the Hypre framework 8_henson_boomeramg_2002 constructs the multilevel hierarchy. The solution methods follow the strategies detailed in Example 1.
Setup for the solution methods: We use the symmetric Successive Over-Relaxation (SSOR) method as a smoother with $3$ pre- and $3$ post-smoothing steps for the multiplicative MG while $6$ smoothing steps on each level for the additive MG, ensuring the same amount of work is done at each level for both multigrid variants. Furthermore, for the MPCG method, we store the last $5$ pairs of vectors to estimate the value of the parameter $\beta$, balancing accuracy and memory use for computational efficiency. While storing more vectors can improve convergence by preserving the conjugacy of the search directions, it becomes memory-intensive, requiring storage proportional to ${(L+1)\times n\times m}$, increasing $m$ enhances convergence but raises memory costs. We store the last $5$ pairs of vectors, as it was shown in4_bridson_multipreconditioned_2006 that a few vectors are sufficient and increasing the number of vectors further might not significantly improve convergence.
@@ 3.1 Comparison with the other solution methods
In this section, we compare the convergence behavior of the additive MG-MPCG method, the additive MG-PCG method, and the multiplicative MG PCG method.
For the Example 1, we consider a high anisotropic case ($k_{xx}=10^{-6}$, $k_{yy}=1$) and a low anisotropic case ($k_{xx}=10^{-1}$, $k_{yy}=1$). As shown in Figure 2, in both cases, the MPCG method outperforms the standard PCG method when preconditioned with the additive MG method. However, as expected, the multiplicative MG-PCG method outperforms both methods. For Example 1, we have employed a multigrid method with geometric hierarchy using the uniform refinement strategy. Hence, we can see that the multilevel hierarchy is not optimal for the highly anisotropic case. Consequently, we observe that the number of iterations required for convergence increases when anisotropy is considered. For the highly anisotropic case, the performance of the additive MG-MPCG method is closer to the MG($V$)-PCG method.
To investigate this behavior further, we carried out the same experiment with more cases with slowly increasing anisotropy. Figure 4 (left) shows that problems with reduced anisotropy converge significantly faster than those with high anisotropy. Among the tested methods, the additive MG-PCG requires the highest number of iterations for convergence, and this requirement amplifies more noticeably with increasing anisotropy. Interestingly, the additive MG-MPCG method, while exhibiting a similar trend, does not experience a significant increase in the number of iterations Its convergence rate grows at a rate comparable to the MG($V$)-PCG methods.
@@
For Example 2, we consider two different test cases with a varying range of fracture permeabilities, where the two extreme cases are chosen, the blocking fractures ($k_{m}=1,k_{f}=10^{-4}$) and the conducting fractures ($k_{m}=1,k_{f}=10^{4}$). In Figure 3, we observe the convergence of the three solution methods for the blocking and conductive fractures. Similar to the previous example, the additive MG-MPCG method outperforms the additive MG-PCG method for both test cases. In this example, we utilize BoomerAMG to build the multilevel hierarchy, which takes into account the permeabilities of the fractures when constructing the transfer operators. Such agglomeration based construction enhances the quality of the multilevel hierarchy compared to that in Example 1, resulting in a more effective hierarchy of the subspaces. We can observe the effect of the transfer operators in Figure 4 (right). In this case, the number of iterations for additive MG-PCG and additive MG-MPCG methods remains relatively constant across different permeabilities, with the additive MG-MPCG method showing a slight increase in iterations as permeability increases.
Optimal transfer operators and smoothers typically ensure a constant number of iterations for additive or multiplicative multilevel methods regardless of problem size. However, our anisotropic problem with non-optimal transfer operators results in increased iterations as problem size grows. Our experiments indicate that the trend observed in Figure 4 will persist with increasing problem size.